# Multi-aspect Knowledge Distillation with Large Language Model

### v3

@@ Abstract
Recent advancements in deep learning have significantly improved performance on computer vision tasks. Previous image classification methods primarily modify model architectures or add features, and they optimize models using cross-entropy loss on class logits. Since they focus on classifying images with considering class labels, these methods may struggle to learn various aspects of classes (e.g., natural positions and shape changes). In contrast, humans classify images by naturally referring to multi-aspects such as context, shape, color, and other features. Inspired by this, rethinking the previous approach from a novel view, we propose a multi-aspect knowledge distillation method using Multimodal Large Language Models (MLLMs). Our approach involves: 1) querying Large Language Model with multi-aspect questions relevant to the knowledge we want to transfer to the model, 2) extracting corresponding logits from MLLM, and 3) expanding the model’s output dimensions to distill these multi-aspect logits. We then apply cross-entropy loss to class logits and binary cross-entropy loss to multi-aspect logits. Through our method, the model can learn not only the knowledge about visual aspects but also the abstract and complex aspects that require a deeper understanding. We primarily apply our method to image classification, and to explore the potential for extending our model, we expand it to other tasks, such as object detection. In all experimental results, our method improves the performance of the baselines. Additionally, we analyze the effect of multi-aspect knowledge distillation. These results demonstrate that our method can transfer knowledge about various aspects to the model and the aspect knowledge can enhance model performance in computer vision tasks. This paper demonstrates the great potential of multi-aspect knowledge distillation, and we believe it offers a promising direction for future research in computer vision and beyond. Our github link is https://github.com/taegyeong-lee/MaKD.
@@ 1 Introduction
Recent advancements in deep learning models have led to significant performance improvements in the field of computer vision, including image classification Vaswani et al. (2017); Vasu et al. (2023); Zhu et al. (2023); Novack et al. (2023), object detection Wu et al. (2023); Ma et al. (2023); Wang et al. (2023). In particular, these advancements, primarily focusing on improving model architectures or incorporating additional features, have greatly enhanced performance in image classification. The methods Vaswani et al. (2017); Liu et al. (2021); Zhu et al. (2023); Tan & Le (2019); He et al. (2016) output class logits and use cross-entropy loss to optimize the models.
However, even if the images in a dataset belong to different classes, they can consist of similar features and make the task more challenging Wei et al. (2021); Parkhi et al. (2012); Krause et al. (2013); Fei-Fei et al. (2004); Wah et al. (2011); Cimpoi et al. (2014). For instance, in CUB200 dataset Wah et al. (2011), most classes share the same features that the superclass “bird” has; i.e. beak, two wings, two legs, and so on. This may require not only the class logit but also additional visual features or aspects that require deeper understanding.
How can humans effectively classify fine-grained images? When classifying fine-grained images, humans not only consider the detailed visual aspects of the given image but also take into account abstract and complex aspects that require a more profound understanding Rong et al. (2021). For example, when given a fine-grained image of a bird, humans might think along the lines of “The beak is sharp,” or “There is a river nearby,” combining both detailed visual features and contextual information.
Inspired by this human ability, the question arises: Could the model’s performance improve if we transfer knowledge about various aspects to it? Multi-modal Large Language Models (MLLMs) have also made significant advancements alongside Large Language Models (LLMs). By taking multi-modal inputs, MLLMs Liu et al. (2024b; 2023); Achiam et al. (2023) can understand and effectively represent visual information, enabling tasks such as visual understanding Guo et al. (2023a); Yang et al. (2022); Tsimpoukelli et al. (2021) and image captioning Li et al. (2023); Zhang et al. (2021); Wang et al. (2021). Additionally, since MLLMs can answer abstract or complex questions, unlike image classification modelsVaswani et al. (2017); Liu et al. (2021); Zhu et al. (2023); Tan & Le (2019); He et al. (2016) that output class logits, we can use MLLMs to transfer various knowledge that may help classification to the model.
@@
First, as shown in Figure 1, we generate questions about the aspects the model aims to learn, based on the classes of the dataset, using the LLM. The generated questions represent the aspects that the model aims to learn during training. Secondly, we provide the generated questions to MLLM to obtain the logits of each aspect. Since MLLM can understand visual information and answer abstract questions, the logits of the MLLM may represent knowledge of the diverse aspects about the dataset. Finally, to distill these extracted multi-aspect logits, we simply expand the dimension of the model’s output by adding the number of aspects to the number of classes, and then we optimize the model by applying cross-entropy loss to the class logits and binary cross-entropy loss to the aspect logits.
@@
We conduct experiments on fine-grained and coarse-grained image classification with various neural networks. Our method outperforms the baselines. Additionally, we analyze the impact of aspect knowledge and discuss the correlations between the aspects and performances of the models. Also, to explore the potential for extending our model, we expand it to other tasks, such as object detection and knowledge distillation.
@@ 2 Related Work
Multimodal Large Language Models. Recently, Multimodal Large Language Models (MLLMs) Achiam et al. (2023); Alayrac et al. (2022); Liu et al. (2024b); Yin et al. (2023); Zhang et al. (2024) have shown significant performance improvements in multi-modal problems such as visual question answering and image captioning by leveraging large-scale datasets to learn a joint embedding space where images and their corresponding textual descriptions are closely aligned. GPT-4o Achiam et al. (2023) has the ability to get the context and has a human-like text generation ability, showing strong performance not only in the natural language processing area but also in multi-modal tasks. InternVL Chen et al. (2024) can address both text and image data and shows better performances in various multimodal tasks (such as visual understanding, language generation, and visual QA) while using fewer computing resources compared to other MLLMs. Motivated by this, we apply the rich knowledge of MLLMs to image classification.
Visual tasks with linguistic information. Many studies Berrios et al. (2023); Menon & Vondrick (2022); Pratt et al. (2023); Yan et al. (2023); Salewski et al. (2024); Yang et al. (2023) try to extract linguistic information from a large language model and use it to settle the visual problems. One method Menon & Vondrick (2022) leverages the linguistic knowledge for each visual category from LLM to generate the descriptions and use the descriptions in zero-shot image classification. Another method Yan et al. (2023) creates the concise set of representative visual attributes from LLM by leveraging their learning-to-search method for interpretable visual recognition. While these methods focus on generating attributes for model training, our approach distills knowledge about various aspects, extending the model’s output dimensions.
Figure 1: Multi-aspect question generation and logit extraction. For multi-aspect question generation (a), we generate various aspect questions from the LLM by using the class and prompt as instructions. For logit extraction about multi-aspect questions (b), we input the generated multi-aspect questions along with the image into the MLLM to extract logits and obtain the probabilities corresponding to yes token.
@@ 3.1 Multi-aspect Question Generation from LLM
Our method is illustrated in Figure 1. First, as shown in Figure 1 (a), we create a total of $N$ multi-aspect questions based on the class labels of the dataset using LLM. Then, considering visual, categorical, and environmental aspects, we filter and select $Q$ multi-aspect questions using the LLM. $Q$ is the number of multi-aspect questions we want to transfer to our model. We use GPT-4o with the system prompt, “You are a good question maker.”, and the instructions, “The dataset consists of $C$ classes and $M$ images. The class list is as follows: $[CLASS]$, Generate $N$ feature-specific yes or no questions, focusing on clear and distinct aspects of the objects in the images in the dataset.” and “Select $Q$ of the most relevant and distinct questions from the list, focusing on various key features that distinguish different class in the dataset.”. These generated aspect questions represent the knowledge we aim to transfer to the models based on datasets.
@@ 3.2 Logit Extraction for Multi-aspect Questions
We generate questions about aspects to be transferred to the model from the LLM. As shown in Figure 1 (b), using an MLLM, we input the dataset and the generated multi-aspect questions, prompting it to answer yes or no. We then extract the logits corresponding to yes and no tokens, and apply the softmax function to both the yes and no logits. We use the softmax results of the yes logits as the targets. Let $i$ be the question index, $z_{\text{y}_{i}}$ be the logit for yes for the $i$-th question and $z_{\text{n}_{i}}$ be the logit for no for the $i$-th question respectively. The softmax probability $q_{i}$ is given by:
@@ 3.3 Expansion of Model Output Dimension
Also, we consider the expanded dimension $D$ such that from $1$ to $C$ is the class logit dimension, and from $C+1$ to $D$ is the aspect logit dimension. The multi-aspect logit dimension is used for the distillation of logits representing the multi-aspect questions. We provide the detail figure in the supplementary materials.
3.4 Mutli-aspect Knowledge Distillation Loss
To distill multi-aspect logits, we extend the model outputs by the number of multi-aspect questions $Q$. The class logit dimension of model output is applied with cross-entropy loss, and the aspect logit dimension is applied with binary-cross entropy loss because we use the probability of the yes token extracted from the MLLM as the target. Let $C$ be the number of classes and $Q$ be the number of multi-aspect questions. We expand the model output to $D$. We apply cross-entropy loss to the outputs from $1$ to $C$ for class classification, and binary-cross entropy loss from $C+1$ to $D$ using multi-aspect probability $q$ as the target.
@@ 3.4 Mutli-aspect Knowledge Distillation Loss
where $\hat{y}$ represents the predicted probability, $y$ are the true labels for the classes, $q$ are the targets for the aspects extracted from the MLLM and $\alpha$ is a factor for balancing the losses. The total loss is defined as follow:
@@ 4.1 Implementation Details
Extract logits of answers from MLLM. According to Section 3.1, we extract the probability values of the yes token about multi-aspect from MLLM. We choose InternVL2-8B Chen et al. (2024) as our MLLM because InternVL2-8B can perform inference on a single NVIDIA RTX 3090 and has strong benchmark performance.
Table 1: Accuracy (%) on the fine-grained image test set. We use a total of six datasets (StanfordCars Krause et al. (2013), OxfordPets Parkhi et al. (2012), DTD Cimpoi et al. (2014), 102Flowers Nilsback & Zisserman (2008), CUB200 Wah et al. (2011), and FGVC-Aircraft Maji et al. (2013)). MLLM is InternVL2-8B. Base is the baseline using cross-entropy loss with class labels. We run each experiment three times and report the average results.
(a) StanfordCars
Zero-shot classification
MLLM 14.30
Base Ours Gap
ResNet18 77.53 83.38 +5.85
ResNet34 80.93 84.33 +3.40
MobileNet-V1 82.84 85.43 +2.59
EfficientNet 86.41 88.07 +1.66
(b) OxfordPets
Zero-shot classification
MLLM 49.38
Base Ours Gap
ResNet18 77.07 82.24 +5.17
ResNet34 79.07 82.78 +3.71
MobileNet-V1 78.12 82.75 +4.63
EfficientNet 83.42 85.27 +1.85
(c) DTD
Zero-shot classification
MLLM 49.20
Base Ours Gap
ResNet18 55.73 59.43 +3.70
ResNet34 53.76 59.89 +6.13
MobileNet-V1 57.22 61.44 +4.22
EfficientNet 60.28 62.87 +2.59
(d) 102Flowers
Zero-shot classification
MLLM 26.88
Base Ours Gap
ResNet18 92.32 94.64 +2.32
ResNet34 92.75 94.89 +2.14
MobileNet-V1 94.14 95.56 +1.42
EfficientNet 95.86 96.78 +0.92
(e) CUB200
Zero-shot classification
MLLM 10.27
Base Ours Gap
ResNet18 53.83 60.07 +6.24
ResNet34 56.48 61.93 +5.45
MobileNet-V1 58.85 63.41 +4.56
EfficientNet 66.04 69.32 +3.28
(f) FGVC-Aircraft
Zero-shot classification
MLLM 11.94
Base Ours Gap
ResNet18 71.76 74.33 +2.57
ResNet34 75.56 76.93 +1.37
MobileNet-V1 78.22 80.41 +2.19
EfficientNet 84.16 84.88 +0.72
Table 2: Accuracy (%) on the coarse-grained image test set. MLLM is InternVL2-8B. Base is the baseline using cross-entropy loss with class labels. We run each experiment three times and report the average results.
(a) Caltech101
Zero-shot classification
MLLM 85.52
Base Ours Gap
ResNet18 73.35 75.77 +2.42
ResNet34 75.36 77.56 +2.20
MobileNet-V1 76.64 79.14 +2.50
EfficientNet 80.05 82.17 +2.12
(b) Mini-ImageNet
Zero-shot classification
MLLM 76.38
Base Ours Gap
ResNet18 76.86 77.72 +0.86
ResNet34 77.47 78.65 +1.18
MobileNet-V1 77.50 78.84 +1.34
EfficientNet 73.05 75.07 +2.02
Fine-grained image classification. We use a total of six datasets: StanfordCars Krause et al. (2013), OxfordPets Parkhi et al. (2012), DTD Cimpoi et al. (2014), 102Flowers Nilsback & Zisserman (2008), CUB200 Wah et al. (2011), and FGVC-Aircraft Maji et al. (2013). For fine-grained image classification, we train all models for 240 epochs, with batch size 16. The initial learning rate is 0.01, divided by 10 at the 150th, 180th and 210th epoch. We use SGD optimizer with the momentum of 0.9, and weight decay is set to 5e-4.
Coarse-grained image classification. We additionally apply our method to the Caltech101 Fei-Fei et al. (2004) and Mini-ImageNet Ravi & Larochelle (2016) datasets for coarse-grained image classification. For Caltech101 Fei-Fei et al. (2004), we train all models for 240 epochs, with batch size 16. The initial learning rate is 0.01, divided by 10 at the 150th, 180th, and 210th epoch. For Mini-ImageNet Ravi & Larochelle (2016), we use the same settings following ImageNet setting of prior work Zhao et al. (2022); Guo et al. (2023b).
More implementation details are included in supplementary materials due to the space limit.
@@ 4.2 Experimental Results
Fine-grained image classification. We mainly focus on fine-grained image classification task. Table 1 shows the experimental results on fine-grained datasets Krause et al. (2013); Parkhi et al. (2012); Cimpoi et al. (2014); Nilsback & Zisserman (2008); Wah et al. (2011); Maji et al. (2013). As shown in Table 1, our method demonstrates significant performance improvements for all models on all datasets compared with the model using cross-entropy loss with class labels. For example, on the StanfordCars dataset with ResNet18, our method shows a 5.85% higher performance compared to the baseline. This indicates that our model effectively transfers knowledge regarding aspects and can help models become more effective when dealing with datasets that have fine-grained features (such as subtle differences in visual appearance and patterns).
Coarse-grained image classification. Additionally, we experiment with our approach on coarse-grained datasets. Table 2 shows the experimental results on Caltech101 Fei-Fei et al. (2004) and Mini-ImageNet Ravi & Larochelle (2016). According to Table 2, our model improves the performance of all baselines. These results indicate that our model is also effective in coarse-grained image classification and demonstrate that transferring diverse knowledge to the model can help improve performance in image classification.

### v4

@@ Abstract
Recent advancements in deep learning have significantly improved performance on computer vision tasks. Previous image classification methods primarily modify model architectures or add features, and they optimize models using cross-entropy loss on class logits. Since they focus on classifying images with considering class labels, these methods may struggle to learn various aspects of classes (e.g., natural positions and shape changes). Rethinking the previous approach from a novel view, we propose a multi-aspect knowledge distillation method using Multimodal Large Language Models (MLLMs). Our approach involves: 1) querying Large Language Model with multi-aspect questions relevant to the knowledge we want to transfer to the model, 2) extracting corresponding logits from MLLM, and 3) expanding the model’s output dimensions to distill these multi-aspect logits. We then apply cross-entropy loss to class logits and binary cross-entropy loss to multi-aspect logits. Through our method, the model can learn not only the knowledge about visual aspects but also the abstract and complex aspects that require a deeper understanding. We primarily apply our method to image classification, and to explore the potential for extending our model, such as object detection. In all experimental results, our method improves the performance of the baselines. Additionally, we analyze the effect of multi-aspect knowledge distillation. These results demonstrate that our method can transfer knowledge about various aspects to the model and the aspect knowledge can enhance model performance in computer vision tasks.
@@ 1 Introduction
Recent advancements in deep learning models have led to significant performance improvements in the field of computer vision, including image classification [31, 30, 44, 20], object detection [36, 16, 33]. In particular, these advancements, primarily focusing on improving model architectures or incorporating additional features, have greatly enhanced performance in image classification. The methods [31, 15, 44, 28, 9] output class logits and use cross-entropy loss to optimize the models.
However, even if the images in a dataset belong to different classes, they can consist of similar features and make the task more challenging [35, 22, 10, 6, 32, 5]. For instance, in CUB200 dataset [32], most classes share the same features that the superclass “bird” has; i.e. beak, two wings, two legs, and so on. This can require not only the class logit but also additional visual features or aspects that require deeper understanding.
How can humans effectively classify fine-grained images? When classifying fine-grained images, humans not only consider the detailed visual aspects of the given image but also take into account abstract and complex aspects that require a more profound understanding [26]. For example, when given a fine-grained image of a bird, humans might think along the lines of “The beak is sharp,” or “There is a river nearby,” combining both detailed visual features and contextual information.
Inspired by this human ability, the question arises: Could the model’s performance improve if we transfer knowledge about various aspects to it? Multi-modal Large Language Models (MLLMs) have also made significant advancements alongside Large Language Models (LLMs). By taking multi-modal inputs, MLLMs [13, 1] can understand and effectively represent visual information, enabling tasks such as visual understanding [7, 39, 29] and image captioning [11, 42, 34]. Additionally, since MLLMs can answer abstract or complex questions, unlike image classification models[31, 15, 44, 28, 9] that output class logits, we can use MLLMs to transfer various knowledge that may help classification to the model.
@@
Firstly, as shown in Figure 1, we generate questions about the aspects. The generated questions not only represent the aspects we aim to transfer to the model during training, but also increase the uncertainty of a class, which can help enable visual understanding. Secondly, we provide the generated questions to MLLM to obtain the logits of each aspect. Since MLLM can understand visual information and answer abstract questions, the logits of the MLLM may represent knowledge of the diverse aspects about the dataset. Finally, to distill these extracted multi-aspect logits, we simply expand the dimension of the model’s output by adding the number of aspects to the number of classes, and then we optimize the model by applying cross-entropy loss to the class logits and binary cross-entropy loss to the aspect logits.
@@
We conduct experiments on fine-grained and coarse-grained image classification with various neural networks. Our method outperforms the baselines. Additionally, we analyze the impact of aspect knowledge and discuss the correlations between the aspects and performances of the models. Also, to explore the potential for extending our model, we expand it to other tasks, such as object detection and Knowledge Distillation (KD).
@@ 2 Related Work
Multimodal Large Language Models. Recently, Multimodal Large Language Models (MLLMs) [1, 2, 14, 40, 41] have shown significant performance improvements in multi-modal problems such as visual question answering and image captioning by leveraging large-scale datasets to learn a joint embedding space where images and their corresponding textual descriptions are closely aligned. GPT-4o [1] has the ability to get the context and has a human-like text generation ability, showing strong performance not only in the natural language processing area but also in multi-modal tasks. InternVL [4] can address both text and image data and shows better performances in various multimodal tasks (such as visual understanding, language generation, and visual QA) while using fewer computing resources compared to other MLLMs. Motivated by this, we apply the rich knowledge of MLLMs to image classification.
Visual tasks with linguistic information. Many studies [3, 18, 23, 37, 27, 38, 21] try to extract linguistic information from a large language model and use it to settle the visual problems. One method [18] leverages the linguistic knowledge for each visual category from LLM to generate the descriptions and use the descriptions in zero-shot image classification. Another method [37] creates a concise set of representative visual attributes from LLM by leveraging their learning-to-search method for interpretable visual recognition. While these methods focus on generating attributes for model training, our approach distills knowledge about various aspects, extending the model’s output dimensions.
Figure 1: Multi-aspect question generation and logit extraction. For multi-aspect question generation (a), we generate various aspect questions from the LLM by using the class and prompt as instructions. For logit extraction about multi-aspect questions (b), we input the generated multi-aspect questions along with the image into the MLLM to extract logits and obtain the probabilities corresponding to yes token. For multi-aspect logit distillation (c), we expand the model’s output dimensions and apply both cross-entropy loss and binary cross-entropy loss.
@@ 3.1 Multi-aspect Question Generation from LLM
Our method is illustrated in Figure 1. First, as shown in Figure 1 (a), we create a total of $N$ multi-aspect questions based on the class labels of the dataset using LLM. We then consider visual, categorical, and environmental aspects in the prompt by focusing on distinguishable features such as shape, color, and background of class images in each dataset and filter and select $Q$ multi-aspect questions for rank prompting using the LLM. $Q$ is the number of multi-aspect questions we want to transfer to our model. We use GPT-4o with the system prompt, “You are a good question maker.”, and the instructions, “The dataset consists of $C$ classes and $M$ images. The class list is as follows: $[CLASS]$, Generate $N$ feature-specific Yes or No questions, focusing on clear and distinct aspects of the objects in the images in the dataset.” and “Select $Q$ of the most relevant and distinct questions from the list, focusing on various key features that distinguish different class in the dataset.”. These generated aspect questions represent the knowledge we aim to transfer to the models based on datasets.
@@ 3.2 Logit Extraction for Multi-aspect Questions
We generate questions about aspects to be transferred to the model from the LLM. As shown in Figure 1 (b), using an MLLM, we input each image in the dataset and the generated multi-aspect questions, prompting it to answer yes or no. We then extract the logits corresponding to yes and no tokens, and apply the softmax function to both the yes and no logits. We use the softmax results of the yes logits as the targets. Let $i$ be the question index, $z_{\text{y}_{i}}$ be the logit for yes for the $i$-th question and $z_{\text{n}_{i}}$ be the logit for no for the $i$-th question respectively. The softmax probability $q_{i}$ is given by:
@@ 3.3 Expansion of Model Output Dimension
StanfordCars OxfordPets DTD 102Flowers
MLLM 14.30 49.38 49.20 26.88
Model Base Ours Gap Base Ours Gap Base Ours Gap Base Ours Gap
ResNet18 77.53 83.38 +5.85 77.07 82.24 +5.17 55.73 59.43 +3.70 92.32 94.64 +2.32
ResNet34 80.93 84.33 +3.40 79.07 82.78 +3.71 53.76 59.89 +6.13 92.75 94.89 +2.14
MobileNet-V1 82.84 85.43 +2.59 78.12 82.75 +4.63 57.22 61.44 +4.22 94.14 95.56 +1.42
EfficientNet 86.41 88.07 +1.66 83.42 85.27 +1.85 60.28 62.87 +2.59 95.86 96.78 +0.92
CUB200 FGVC-Aircraft Caltech101 Mini-ImageNet
MLLM 10.27 11.94 85.52 76.38
Model Base Ours Gap Base Ours Gap Base Ours Gap Base Ours Gap
ResNet18 53.83 60.07 +6.24 71.76 74.33 +2.57 73.35 75.77 +2.42 76.86 77.72 +0.86
ResNet34 56.48 61.93 +5.45 75.56 76.93 +1.37 75.36 77.56 +2.20 77.47 78.65 +1.18
MobileNet-V1 58.85 63.41 +4.56 78.22 80.41 +2.19 76.64 79.14 +2.50 77.50 78.84 +1.34
EfficientNet 66.04 69.32 +3.28 84.16 84.88 +0.72 80.05 82.17 +2.12 73.05 75.07 +2.02
Table 1: Accuracy (%) on the fine-grained and coarse-grained image test set. For fine-grained image classification task, we use a total of 6 datasets (StanfordCars [10], OxfordPets [22], DTD [5], 102Flowers [19], CUB200 [32], and FGVC-Aircraft [17]). For coarse-grained image classification task, we use a total of 2 datasets (Caltech101 [6] and Mini-ImageNet [24]). MLLM is InternVL2-8B. Base is the baseline using cross-entropy loss with class labels. We run each experiment three times and report the average results.
Also, we consider the expanded dimension $D$ such that from $1$ to $C$ is the class logit dimension, and from $C+1$ to $D$ is the aspect logit dimension. The multi-aspect logit dimension is used for the distillation of logits representing the multi-aspect questions.
3.4 Multi-aspect Knowledge Distillation Loss
As shown in Figure 1, to distill multi-aspect logits, we extend the model outputs by the number of multi-aspect questions $Q$. The class logit dimension of model output is applied with cross-entropy loss, and the aspect logit dimension is applied with binary-cross entropy loss because we use the probability of the yes token extracted from the MLLM as the target. We apply cross-entropy loss to the outputs from $1$ to $C$ for class classification, and binary-cross entropy loss from $C+1$ to $D$ using multi-aspect probability $q$ as the target. We refer to the binary cross-entropy loss for multi-aspect questions as the multi-aspect knowledge distillation (MaKD) loss.
@@ 3.4 Multi-aspect Knowledge Distillation Loss
Where $\hat{y}$ represents the predicted probability from MLLM, $y$ are the true labels for the classes, $q$ are the targets for the aspects extracted from the MLLM and $\alpha$ is a factor for balancing the losses. The total loss is defined as follow:
@@ 4.1 Implementation Details
Extract logits of answers from MLLM. According to Section 3.1, we extract the probability values of the yes token about multi-aspect from MLLM. We choose InternVL2-8B [4] as our MLLM because InternVL2-8B can perform inference on a single NVIDIA RTX 3090 and has strong benchmark performance.
Fine-grained image classification. We use a total of six datasets: StanfordCars [10], OxfordPets [22], DTD [5], 102Flowers [19], CUB200 [32], and FGVC-Aircraft [17]. For fine-grained image classification, we train all models for 240 epochs, with batch size 16. The initial learning rate is 0.01, divided by 10 at the 150th, 180th and 210th epoch. We use SGD optimizer with the momentum of 0.9, and weight decay is set to 5e-4.
Coarse-grained image classification. We additionally apply our method to the Caltech101 [6] and Mini-ImageNet [24] datasets for coarse-grained image classification. For Caltech101 [6], we train all models for 240 epochs, with batch size 16. The initial learning rate is 0.01, divided by 10 at the 150th, 180th, and 210th epoch. For Mini-ImageNet [24], we use the same settings following ImageNet setting of prior work [43, 8].
@@ 4.2 Experimental Results
Fine-grained image classification. We mainly focus on fine-grained image classification task. Table 1 shows the experimental results on fine-grained datasets [10, 22, 5, 19, 32, 17]. As shown in Table 1, our method demonstrates significant performance improvements for all models on all datasets compared with the model using cross-entropy loss with class labels. For example, on the StanfordCars dataset with ResNet18, our method shows a 5.85% higher performance compared to the baseline. This indicates that our model effectively transfers knowledge regarding aspects and can help models become more effective when dealing with datasets that have fine-grained features (such as subtle differences in visual appearance and patterns).
Coarse-grained image classification. Additionally, we experiment with our approach on coarse-grained datasets. Table 1 shows the experimental results on Caltech101 [6] and Mini-ImageNet [24]. According to Table 1, our model improves the performance of all baselines. These results indicate that our model is also effective in coarse-grained image classification and demonstrate that transferring diverse knowledge to the model can help improve performance in image classification.